# Temporal Action Localization for Inertial-based Human Activity Recognition

### v1

@@ 5. Results
Table 2 provides average results of the seven tested architectures across each dataset. One can see that the TAL architectures outperform the inertial architectures across all datasets regarding average mAP. This shows that by being trained to specifically optimize activity boundaries, the different prediction target has resulted in overall more coherent segments which overlap to a larger degree with the ground truth segments. Even though prediction results of the inertial architectures were further smoothed by a majority filter, average mAP is, except for the WEAR dataset, more than halved when compared to that of the TAL architectures. Regarding traditional classification metrics the TAL architectures are able to outperform inertial architectures for four out of six datasets with only the RWHAR (Sztyler and Stuckenschmidt, 2016) and WEAR dataset (Bock et al., 2023) being the exception. These improvements range between $5\%$ for the Opportunity and Wetlab dataset, $10\%$ for the Hang-Time and even as much as 25% in F1-score for the SBHAR dataset. Calculated misalignment ratios show that both inertial and TAL architectures have a similar distribution of errors, with architectures producing better overall classification results also showing overall lower misalignment measures. Only for the RWHAR dataset one can see that TAL architectures show a significantly higher Overfill-Ratio. This might be due to the RWHAR dataset not featuring a NULL-class, which introduces an uncommon prediction scenario for the TAL architectures. Nevertheless, the performed DETAD analysis (Alwassel et al., 2018), which further differentiates amongst the segment-based errors, reveals that inertial architectures suffer more severely from background errors, i.e. confusing activities with the NULL-class. While this effect is decreased for the shallow DeepConvLSTM, TAL architectures show to be able to more reliably differentiate between activities and the NULL-class, and thus more reliably localize activities within the untrimmed sequences.
@@
Table 2. Average LOSO results obtained on six inertial HAR benchmark datasets (Roggen et al., 2010; Reyes-Ortiz et al., 2016; Scholl et al., 2015; Bock et al., 2023; Hoelzemann et al., 2023; Sztyler and Stuckenschmidt, 2016) for four inertial (Ordóñez and Roggen, 2016; Bock et al., 2021; Abedin et al., 2021; Zhou et al., 2022) and three TAL architectures (Zhang et al., 2022b; Tang et al., 2023; Shi et al., 2023). The table provides per-sample classification metrics, i.e. Precision (P), Recall (R), F1-Score (F1), misalignment ratios (Ward et al., 2006) and average mAP applied at different tIoU thresholds (0.3:0.1:07). All experiments employed a sliding window of one second with a $50\%$ overlap. Results are averaged across three runs employing different random seeds. The TAL architectures are able to outperform the inertial architectures regarding average mAP on all HAR datasets and result in the highest classification metrics on four out of the six datasets. Best results per dataset are in bold.
@@
Per-Class Results
Figures 6 and 7 show the confusion matrices of the TAL and inertial architectures obtained on the SBHAR and RWHAR datasets. With the improvements on the SBHAR being the largest across all datasets, one can see that this increase can be attributed to improved performance on transitional, non-periodic classes like sit-to-stand. These activities are mostly recognisable by their context and surrounding activities and are thus particularly challenging to predict for models that do not rely on temporal dependencies spanning multiple seconds. Since the DeepConvLSTM, Attend-and-Discriminate and TinyHAR models are being trained on shuffled training data and have recurrent parts applied on within-window sequences, the three architectures cannot learn to predict these activities based on surrounding context. Among inertial models, only the shallow DeepConvLSTM is able to more reliably predict these context-based classes as it is trained on unshuffeled data and applies a dimensional flip when training the LSTM. Across all datasets part of this benchmark analysis, the RWHAR dataset yields the least performant results for the TAL architectures. We accredit this primarily due to the absence of a NULL-class in the dataset, which introduces a uncommon prediction scenario for the models. Furthermore, the RWHAR consists of the least amount of segments per subject, limiting the amount of training segments which can be used to optimize the TAL models. Results on the RWHAR dataset are nevertheless surprising as the TAL models show confusion among classes which they do not struggle to predict in other datasets (e.g., lying in the SBHAR dataset) as well as classes that are not similar to each other (e.g., lying and jumping). The obtained results raise the question of whether TAL models are primarily suited for being applied on untrimmed sequences, which (1) include breaks and/ or (2) provide a larger amount of segments than the RWHAR dataset.
@@
Figure 6. Confusion matrices of the ActionFormer (Zhang et al., 2022b), TemporalMaxer (Tang et al., 2023) and TriDet model (Shi et al., 2023) applied on the SBHAR (Reyes-Ortiz et al., 2016) (top) and RWHAR dataset (Sztyler and Stuckenschmidt, 2016) (bottom) with a one second sliding window and 50% overlap. Results are postprocessed using a score threshold of 0.3 in case of the SBHAR dataset. Note that confusions which are 0 are omitted.
@@
Figure 7. Confusion matrices of the DeepConvLSTM (Ordóñez and Roggen, 2016), shallow DeepConvLSTM (Bock et al., 2021), Attend-and-Discriminate (Abedin et al., 2021) and TinyHAR model (Zhou et al., 2022) applied on the SBHAR (Reyes-Ortiz et al., 2016) (top) and RWHAR dataset (Sztyler and Stuckenschmidt, 2016) (bottom) with a one second sliding window and 50% overlap. Results are postprocessed using a majority filter 5 seconds in case of the SBHAR dataset. Note that confusions which are 0 are omitted.
@@ 6. Discussion & Conclusion
This article demonstrated the applicability of vision-based, single-stage Temporal Action Localization for inertial-based Human Activity Recognition. We showed that three state-of-the-art TAL models (Zhang et al., 2022b; Tang et al., 2023; Shi et al., 2023) can be applied in a plain fashion to raw inertial data and achieve competitive results on six popular inertial HAR datasets (Roggen et al., 2010; Reyes-Ortiz et al., 2016; Scholl et al., 2015; Sztyler and Stuckenschmidt, 2016; Hoelzemann et al., 2023; Bock et al., 2023), outperforming in most cases popular models from the inertial community by a significant margin. A previously unexplored metric in inertial-based HAR, mean Average Precision (mAP), reveals that TAL models predict less fragmented timelines compared to inertial models and overall achieve larger degrees of overlap with ground truth segments. Furthermore, TAL models show to predict even non-periodic and complex activities more reliably than inertial architectures, providing consistent results across all types of classes across all datasets. Being one of the key challenges in HAR (Bulling et al., 2014), the TAL architectures further show to be less affected by the unbalanced nature of HAR datasets due to a large NULL-class. Across the five benchmark datasets which offered a NULL-class, the TAL architectures showed to deliver the highest NULL-class accuracy.
Within the past years inertial-based HAR has been working on improving upon methods to model temporal relationships, such as using Transformers as opposed to LSTMs to handle longer sequences (Zhou et al., 2022). Nevertheless, as also evident from this benchmark analysis, the way these mechanisms are used in popular inertial models does not give them the capability to learn temporal context beyond the current timestamp which is to be predicted (Pellatt and Roggen, 2020). This causes predicted timelines to suffer from frequent activity switches. Even though researchers have suggested new metrics (Ward et al., 2011), traditional classification metrics remain the focus of analyses within the inertial community. These metrics, though providing a good intuition about the overall recognition performance, do not provide any insight about the degree of fragmentation of the predicted timeline. With researchers such as Pellatt and Roggen (2020) and Guan and Plötz (2017) having worked on implementing improved training strategies, inertial architectures by default still lack to incorporate context-aware information into their models. Nevertheless, results obtained using the shallow DeepConvLSTM demonstrate that inertial models can be set to learn temporal context, yet require more segment-based optimization to be comparable to TAL architectures as far as average mAP scores.
To summarize, the way TAL architectures are trained does not allow them to be applied in an online fashion. Furthermore, comparing learnable parameters of TAL and inertial models, the former is substantially larger and thus is not suited to be deployed on edge devices. Nevertheless, the retrospective analysis of a fully-captured timeline performed by the TAL community along with fast training times even on large sequences offers an interesting new perspective to inertial-based HAR. Results show that these types of models are leveraging context-based information and are able to outperform popular inertial models on a variety of datasets. Even though models of the two independent communities are sharing similarities, the TAL community offers many unexplored design choices and training concepts, which should be considered to be investigated by the inertial-based HAR community.

### v2

@@ 5.1. Offline Activity Recognition
Table 2 provides average results of the seven tested architectures across each dataset in an offline prediction setting. One can see that the TAL architectures outperform the inertial architectures across all datasets regarding average mAP. This shows that by being trained to specifically optimize activity boundaries, the different prediction target has resulted in overall more coherent segments which overlap to a larger degree with the ground truth segments. Even though prediction results of the inertial architectures were further smoothed by a majority filter, average mAP is, except for the WEAR dataset, more than halved when compared to that of the TAL architectures. Regarding traditional classification metrics the TAL architectures are able to outperform inertial architectures for four out of six datasets with only the RWHAR (Sztyler and Stuckenschmidt, 2016) and WEAR dataset (Bock et al., 2023) being the exception. These improvements range between $5\%$ for the Opportunity and Wetlab dataset, $10\%$ for the Hang-Time and even as much as 25% in F1-score for the SBHAR dataset. Calculated misalignment ratios show that both inertial and TAL architectures have a similar distribution of errors, with architectures producing better overall classification results also showing overall lower misalignment measures. Only for the RWHAR dataset one can see that TAL architectures show a significantly higher Overfill-Ratio. This might be due to the RWHAR dataset not featuring a NULL-class, which introduces an uncommon prediction scenario for the TAL architectures. Nevertheless, the performed DETAD analysis (Alwassel et al., 2018), which further differentiates amongst the segment-based errors, reveals that inertial architectures suffer more severely from background errors, i.e. confusing activities with the NULL-class. While this effect is decreased for the shallow DeepConvLSTM, TAL architectures show to be able to more reliably differentiate between activities and the NULL-class, and thus more reliably localize activities within the untrimmed sequences.
@@
Table 2. Offline Activity Recognition: Average LOSO cross-validation results obtained on six inertial HAR benchmark datasets (Roggen et al., 2010; Reyes-Ortiz et al., 2016; Scholl et al., 2015; Bock et al., 2023; Hoelzemann et al., 2023; Sztyler and Stuckenschmidt, 2016) for four inertial (Ordóñez and Roggen, 2016; Bock et al., 2021; Abedin et al., 2021; Zhou et al., 2022) and three TAL architectures (Zhang et al., 2022b; Tang et al., 2023; Shi et al., 2023). The table provides per-sample classification metrics, i.e. Precision (P), Recall (R), F1-Score (F1), misalignment ratios (Ward et al., 2006) and average mAP applied at different tIoU thresholds (0.3:0.1:07). All experiments employed a sliding window of one second with a $50\%$ overlap. Results are averaged across three runs employing different random seeds. The TAL architectures are able to outperform the inertial architectures regarding average mAP on all HAR datasets and result in the highest classification metrics on four out of the six datasets. Best results per dataset are in bold.
@@
With the improvements on the SBHAR being the largest across all datasets, one can see on a per class level (see Figure 5) that this increase can be attributed to improved performance on transitional, non-periodic classes like sit-to-stand. These activities are mostly recognisable by their context and surrounding activities and are thus particularly challenging to predict for models that do not rely on temporal dependencies spanning multiple seconds. By applying the TAL architectures in an offline manner, the architectures are able to leverage both local and global context across the whole timeline and are thus even able to recognize these short-lasting activities. Since the DeepConvLSTM, Attend-and-Discriminate and TinyHAR models are being trained on shuffled training data and have recurrent parts applied on within-window sequences, the three architectures cannot learn to predict these activities based on surrounding context. Among inertial models, only the shallow DeepConvLSTM is able to more reliably predict these context-based classes as it is trained on unshuffeled data and applies a dimensional flip when training the LSTM. Across all datasets part of this benchmark analysis, the RWHAR dataset yields the least performant results for the TAL architectures. We accredit this primarily due to the absence of a NULL-class in the dataset, which introduces a uncommon prediction scenario for the models. Furthermore, the RWHAR consists of the least amount of segments per subject, limiting the amount of training segments which can be used to optimize the TAL models. Results on the RWHAR dataset are nevertheless surprising as the TAL models show confusion among classes which they do not struggle to predict in other datasets (e.g., lying in the SBHAR dataset) as well as classes that are not similar to each other (e.g., lying and jumping). The obtained results raise the question of whether TAL models are primarily suited for being applied on untrimmed sequences, which (1) include breaks and/ or (2) provide a larger amount of segments than the RWHAR dataset.
@@
Figure 5. Offline Activity Recognition: Confusion matrices of the (a) the best TAL architecture (TriDet) (Shi et al., 2023) and (b) inertial model (shallow DeepConvLSTM) being applied on the SBHAR (Reyes-Ortiz et al., 2016) (top) and RWHAR dataset (Sztyler and Stuckenschmidt, 2016) (bottom) with a one second sliding window and 50% overlap. Note that confusions which are 0 are omitted.
[…]
@@
Offline Activity Recognition: Confusion matrices of the (a) the best TAL architecture (TriDet) and (b) inertial model (shallow DeepConvLSTM) being applied on the SBHAR (Reyes-Ortiz et al., 2016) (top) and RWHAR dataset (bottom) with a one second sliding window and 50% overlap. Note that confusions which are 0 are omitted.
Single- vs. Two-Stage TAL Training
As described in Section 3, the TAL models described in this paper were intended to be applied using feature embeddings describing each sliding window, rather than raw data. Nevertheless, our initial results show that TAL models are indeed capable of being applied to raw inertial data. Figure 6 presents results of our implemented two-stage TAL training, which extends the training process as described in the previous chapters with a prepended feature extraction using inertial models. In total, we assess two feature embeddings: LSTM-based features extracted from a DeepConvLSTM (Ordóñez and Roggen, 2016) and attention-based features from a TinyHAR architecture (Zhou et al., 2022). Though the two-stage training using LSTM-based features only yields better F1-scores for the Wetlab, WEAR, and RWHAR datasets, the training using attention-based features improves results across all datasets across all TAL models. Given the significant improvements on the RWHAR and WEAR datasets, we assume that the feature extraction via the prepended inertial network helps increase the discriminability of the window-level features, yet at the cost of making it harder for the TAL models to learn cross-window temporal relations, as evident by the decrease in mAP scores across (almost) all datasets.
Figure 6. Offline Activity Recognition: Average LOSO results of the two-stage TAL training of the TriDet model. We compare the LSTM and attention feature-based two-stage TAL training with the best inertial-based architecture and the single-stage TAL training on raw inertial data of all six benchmark datasets. One can see a clear improvement when using attention-based features across all datasets.
[…]
Offline Activity Recognition: Average LOSO results of the two-stage TAL training of the TriDet model. We compare the LSTM and attention feature-based two-stage TAL training with the best inertial-based architecture and the single-stage TAL training on raw inertial data of all six benchmark datasets. Results are presented as bar plots. One can see a clear improvement when using attention-based features across all datasets.
5.2. Online Activity Recognition
Table 3. Average LOSO results obtained using the evaluated TAL architectures (Zhang et al., 2022b; Tang et al., 2023; Shi et al., 2023) employing different degrees of chunking during validation. During validation each subject is split into equal-sized chunks with each chunk being individually predicted by the trained TAL models. The table provides per-sample F1-Score (F1), average mAP applied at different tIoU thresholds (0.3:0.1:07) calculated on each chunk individually (c-mAP) and on the reconstructed validation stream (mAP). Results are averaged across three runs employing different random seeds. Results which are underlined are not score thresholded.
1 sec. 5 sec. 30 sec. 60 sec. Unchunked
Model F1 c-mAP mAP F1 c-mAP mAP F1 c-mAP mAP F1 c-mAP mAP F1 mAP
Opp. AF 2.63 2.45 0.36 37.73 18.10 25.51 49.96 41.74 45.84 50.50 45.45 47.84 51.93 51.24
TM 2.22 2.09 0.31 38.00 19.08 25.99 45.32 39.24 42.83 45.27 42.15 44.22 44.74 46.31
TD 1.48 2.26 0.37 9.96 10.51 13.6 12.15 12.73 13.37 1.99 4.15 4.27 48.79 49.70
SBHAR AF 3.44 7.56 0.12 34.49 26.20 15.09 78.88 74.28 79.52 81.57 82.45 86.39 84.43 95.46
TM 3.30 5.39 0.05 33.49 47.11 23.87 78.08 73.95 77.64 81.04 82.41 85.19 83.66 94.39
TD 2.13 10.18 0.13 17.69 29.31 8.05 31.70 52.01 49.38 44.76 72.94 72.64 86.45 94.75
Wetlab AF 5.13 2.37 0.02 14.55 3.12 0.77 35.35 25.65 20.15 35.98 30.46 28.06 40.71 33.53
TM 4.82 2.23 0.03 11.33 15.81 1.51 34.37 25.87 20.16 35.35 30.24 26.90 37.09 35.72
TD 1.77 4.08 0.03 2.20 4.40 0.63 4.03 6.10 7.38 7.78 10.98 9.11 42.85 34.05
WEAR AF 5.70 3.47 0.00 15.54 10.44 0.06 53.57 40.02 18.72 69.21 56.89 52.11 72.43 73.80
TM 5.30 2.65 0.00 38.37 53.39 0.19 58.29 45.09 18.94 67.91 56.60 47.89 69.52 69.18
TD 1.25 5.41 0.00 3.72 15.12 0.09 34.67 49.73 21.57 43.96 49.33 43.79 73.18 77.12
Hang-T. AF 6.41 3.19 0.70 42.41 17.18 14.67 49.32 26.47 26.82 50.69 27.64 27.65 51.23 29.26
TM 6.39 2.63 0.50 41.13 17.54 15.17 48.15 25.72 25.38 48.67 26.63 26.33 47.17 27.86
TD 19.52 0.00 3.56 19.33 3.67 3.48 26.02 12.42 13.01 35.72 18.80 19.15 50.67 29.24
RWHAR AF 8.47 28.67 0.00 17.00 51.96 0.00 50.15 76.19 1.37 62.40 77.58 7.91 61.24 65.40
TM 8.06 17.13 0.00 31.35 48.89 0.00 52.71 60.47 1.31 58.47 61.86 7.61 60.59 50.60
TD 5.88 25.68 0.00 9.07 40.11 0.00 39.78 47.78 0.94 48.65 57.51 7.39 67.86 69.98
A major difference in how TAL models are intended to be used compared to inertial models is that they can only provide an offline prediction of previously recorded timeseries data. In previous experiments, we demonstrated that TAL models are capable of analyzing a participant’s data stream as a whole and can even deal with arbitrary lengths of timeseries data and activity durations. However, to investigate how capable TAL models are at analyzing only small snippets of incoming data streams, we modified the previously employed Leave-One-Subject-Out validation loop so that TAL models are tasked to predict each validation subjects’ data in a chunked manner. Specifically, given the data of a validation participant, we divided the data into equal-sized portions (e.g. 5 seconds worth of data) and had the TAL models, trained using unchunked data, predict each chunk individually. Table 3 compares the initial unchunked results of the tested TAL models (ActionFormer, TemporalMaxer and TriDet) on the six inertial benchmark datasets with results obtained when fragmenting each validation split into 1, 5, 30, or 60-second chunks. Across all datasets one can witness that with smaller-sized chunks classification and mAP results worsen across all algorithms and datasets. Overall, we can see that the smaller the chunk, the lower the overall confidence of the TAL models regarding each predicted segment. The lower confidence further leds to almost all segment predictions being removed even with a low score threshold such as 0.05, resulting in us not applying score thresholding in those cases, i.e., using all segments predicted by the respective TAL model. We assume that this drop in prediction confidence is likely caused by artificially splitting long segments of activities, contradicting what the models have seen during training, as datasets such as the Opportunity and Hang-Time dataset, which contain mostly short-lasting segments, are less affected by the smaller chunks as e.g. the WEAR dataset which has almost exclusively segments lasting longer than 30 seconds. While the TriDet model is not capable of dealing with the chunked validation, most likely due to it using a different style of regression head, the ActionFormer and the TemporalMaxer are capable of maintaining predictive performance even for small-sized chunks of 5 seconds across almost all datasets. As expected, mAP scores are more impacted by the chunked prediction as chances are increased that long lasting segments are split due to the predicted segments not spanning across chunks. Nevertheless, we expect this effect to dampen if one would use an additional majority filtering as used in our inertial-based experiments, as this would eliminate potential intermediate activity switches.
Table 4 provides a comparison in terms of learnable parameters, size on disk, average batch training and inference times of the inertial-based models compared with the TAL models. One can see that the TAL models are significantly larger in size and number of learnable parameters. Though all architectures part of our analysis are capable of predicting a window of one second as well as the complete prediction stream of a participant within less than half a second, inertial-based models are faster than TAL models in both inference scenarios being on average faster by around a magnitude of 10. Taking into consideration that all TAL models fail to predict 1 second chunks and are significantly larger than inertial-based models, it becomes aparent, that TAL models are not suited for inference on edge-devices. It becomes apparent that with the models being trained to predict activity occurrences as a whole, the models require enough surrounding context in order to spot activities and correctly identify what they have seen during training. Nevertheless, as soon as the TAL models are given enough surrounding context (e.g., a prior context of 60 seconds), the models are capable to spot activities within said activity stream in a reasonable run time.
Table 4. Comparison in terms of learnable parameters (in million (M)), size on disk (in MB), average batch training and inference times (in milliseconds) of the seven benchmark algorithms. Training and inference speeds are the average across the first 5 epochs of the first LOSO validation split of the algorithms being applied on the WEAR dataset. Benchmarking was performed on a single NVIDIA GeForce RTX 4090 with an AMD Ryzen 7800 X3D. We assess both inference speeds of the architectures being tasked to predict one 1-second window as well as one complete subject. ∗∗ 1 batch equals 100 1-second windows ∗∗ 1 batch equals 1 participant.
Model Total Params Size Avg. Time per Batch
Train Test (1 sec.) Test (1 participant)
Baseline DeepConvLSTM 0.71 M 2.69MB 4ms 2ms 27ms
Shallow D. 0.57 M 2.19MB 3ms 2ms 29ms
A-and-D 0.57 M 2.19MB 31ms 10ms 41ms
TinyHAR 0.04 M 0.14MB 7ms 3ms 28ms
ActionFormer 27.02 M 103.37MB 131ms∗∗ 39ms 346ms
TemporalMaxer 4.89 M 18.94MB 69ms∗∗ 25ms 380ms
TriDet 13.75 M 52.73MB 127ms∗∗ 37ms 347ms
Figure 7. Color-coded visualization of the effect of the Offline and Online Activity Recognition experiments using the ActionFormer model (Zhang et al., 2022b). Results are compared with the ground truth of a sample participant as well as offline single-stage training for both the RWHAR (Sztyler and Stuckenschmidt, 2016) and SBHAR (Reyes-Ortiz et al., 2016) dataset. One can see that training the TAL models using attention based features clearly improves results with an overall better recognition of all activities. Contrarily, once predicting a chunked output, performance can be somewhat maintained for larger chunks, yet drops near zero once applied on only 5-second long windows.
[…]
Visualization of the prediction streams of the ActionFormer for single-stage, two-stage and chunked prediction scenarios. Results are visualized for both the SBHAR and RWHAR dataset and compared against ground truth of a sample participant.
@@ 6. Discussion & Conclusion
This article demonstrated the applicability of vision-based, single-stage Temporal Action Localization for inertial-based Human Activity Recognition. We showed that three state-of-the-art TAL models (Zhang et al., 2022b; Tang et al., 2023; Shi et al., 2023) can be applied in a plain fashion to raw inertial data and achieve competitive results on six popular inertial HAR datasets (Roggen et al., 2010; Reyes-Ortiz et al., 2016; Scholl et al., 2015; Sztyler and Stuckenschmidt, 2016; Hoelzemann et al., 2023; Bock et al., 2023), outperforming in most cases popular models from the inertial community in an offline prediction scenario by a significant margin. Using a combination of an inertial network as a feature extractor and a TAL model showed to significantly enhance classification results for all TAL models – especially on datasets where classic inertial models had previously outperformed TAL models. Our two-stage experiments further suggest potential improvements that could be achieved by investigating improved methods for feature extraction using TAL models and inertial data, e.g. via a fully-differentiable combined version of both type of architectures (see Figure 7).
@@
A previously unexplored metric in inertial-based HAR, mean Average Precision (mAP), reveals that TAL models predict less fragmented timelines compared to inertial models and overall achieve larger degrees of overlap with ground truth segments. Furthermore, TAL models show to predict even non-periodic and complex activities more reliably than inertial architectures, providing consistent results across all types of classes across all datasets. Being one of the key challenges in HAR (Bulling et al., 2014), the TAL architectures are further shown to be less affected by the unbalanced nature of HAR datasets due to a large NULL-class. Across the five benchmark datasets which offered a NULL-class, the TAL architectures showed to deliver the highest NULL-class accuracy. Additional experiments, which involved applying the TAL models on artificially created chunked sequences of data, showed that TAL models, though intended for the offline analysis of prerecorded activity timelines, have the capability of being applied in a near-online fashion (see Figure 7). Although the TAL models highlighted in this paper overall size and complexity would not allow them to be run on edge devices (yet), their reasonable performance and inference time on large enough chunks suggest that they could function e.g. as a server-side prediction model as seen in (Berchtold et al., 2010), which analyzes chunks of data in regular intervals.
The research community for inertial-based activity recognition has contributed methods to better model temporal relationships in the past years, yet most such architectures are limited to learning context within a fixed-sized sliding window. To this date, the length of the sliding window remains a crucial parameter in inertial-based HAR, which may result in a significant performance drops in recognition performance when set incorrectly. Specifically, if dealing with both long and short lasting activities, a small sliding window might end up being too small to fully capture an activity, while a too large window size might cause sliding windows containing mixed activity types. With studies such as Pellatt and Roggen (2020) and Guan and Plötz (2017) contributing strategies to improve training, the TAL community offers an interesting new perspective to inertial-based HAR – alleviating the need for fixed size windows and making algorithms capable of dealing with activities of arbitrary length. Although models from these two independent communities share similarities, the TAL community offers many unexplored design choices and training concepts for a multitude of application scenarios, which we argue should be considered for investigation by the inertial-based HAR community.
Acknowledgements.
We gratefully acknowledge the DFG Project WASEDO (grant number 506589320) and the University of Siegen’s OMNI cluster.